# entity resolution AL active learning

### v2

@@ 5.2. Experimental Setup
MultiEM (Zeng et al., 2024). As proposed in the original work (Zeng et al., 2024), we apply a grid search to identify the best configuration for each dataset and report the best result based on the unsolved entity resolution problems. Due to the involvement of data sources with duplicates, we also test $k\geq 1$ to determine candidates for each record.
@@
Ditto (Li et al., 2020). For faster training, we utilize DistilBERT (Sanh et al., 2019) as a pre-trained model. Moreover, we use a maximum sequence length of 256, also fixed in the original work (Li et al., 2020), and train the model for 10 epochs. We report the results that lead to the smallest training loss over the 10 epochs.
We evaluate AnyMatch (Zhang et al., 2025) using its official GitHub implementation [^6] https://github.com/Jantory/anymatch. For each dataset, the model is trained on the corresponding training set and evaluated on the test set, with the parameterized sample size $n_{r}$ for comparability. Note, AnyMatch filters relevant record pairs using the full ground truth of the training data. We use a batch size of 32, GPT-2 as the language model, and a learning rate of 2e-5.
Table 4. Summary of speedup factors for MoRER with different AL methods compared to the approaches Almser (Alm), TransER (TER), Sudowoodo (Su), Ditto (Dit), and AnyMatch (Any) regarding the datasets Dexter (Dex), Music (Mus), and WDC-computer (WDC).
@@
Dex 27.3 102.7 127.3 46 410.5 413.8 132.2 28.1 78.0 96.6 41.2 311.6 314.1 106.1 28.9 58.7 72.8 33.4 234.6 236.5 76.8
Mus 75.7 40.6 33.1 80.6 115.4 87.4 74.3 21.9 24.8 20.2 49.9 70.6 53.5 14 63.8 18.8 15.3 40 53.4 40.4 11.2
WDC 9.8 2.4 2.3 306 21.0 15.7 17.3 13.2 1.9 1.8 240.5 16.4 12.2 15.4 15.0 1.4 1.3 176.7 11.8 8.8 11.2
5.3. Comparison to Baselines
Effectiveness
The results, shown in Fig. 5, indicate that MoRER performs comparably to Almser under certain budgets and active learning configurations. MoRER, when combined with Almser, outperformed both the Bootstrap method and Almser alone on the Dexter dataset. We hypothesize that Almser benefits from the high number of entity resolution tasks, resulting in rich similarity graphs, even when these graphs are divided due to the separation of entity resolution tasks. Moreover, the partitioning of entity resolution tasks reduces the search space for identifying informative links for similar classes of entity resolution tasks, thereby minimizing the likelihood of selecting uninformative feature vectors.
In contrast, TransER consistently underperformed relative to the combined performance of Almser and MoRER across all tasks, regardless of whether the full training dataset or the 50% version was utilized. This consistent underperformance shows its limitations in handling large volumes of heterogeneous training data, which is particularly evident in the Dexter and WDC-computer datasets. Considering the language model-based methods without using labeled training data, Sudowoodo and MultiEM achieve comparable results compared to the aforementioned methods according to the Music dataset. However, the approaches result in a decreased F1 score of up to 38% (Dexter) when comparing Sudowoodo with MoRER+Almser, and a decrease of 35% (WDC-computer) when considering MultiEM and MoRER+Almser. Sudowoodo was originally conceptualized for pairwise record linkage. Therefore, we assume that the various and heterogeneous ER problems in Dexter and WDC-computer cause the challenge of determining representations to distinguish matches from non-matches. We hypothesize that the hierarchical processing of MultiEM during the merging phase leads to suboptimal record embeddings, resulting in ineffective classifications.
AnyMatch achieves comparable performance only on the Music dataset. For the WDC-computer dataset, the results decrease by up to 22% compared to MoRER+Almser and by 17.4% when using MoRER with Bootstrap. We assume this decrease is caused by the limitations of the AutoML selection, which appears insufficient for building effective models when classifying large numbers of record pairs.
@@ 5.3. Comparison to Baselines
Efficiency
MoRER combined with the Bootstrap AL method demonstrates significant advantages, outperforming both Almser and TransER across all datasets and labeling budgets, as shown in  Fig. 6. For example, on the Music dataset, MoRER+Bootstrap runtimes range from 121 to 2621 seconds, whereas Almser can take up to 4.6 hours. Almser’s longer runtimes are due to the increasing size of similarity graphs, which require more computational resources for graph analysis. The efficiency advantage results from the MoRER strategy of clustering similar ER tasks among the integrated data sources to select training data. This clustering approach significantly decreases the search space for potential links, contributing to the observed performance improvements. Reducing the search space also positively influences the combination of MoRER and Almser. The performance boost results in a maximum speedup of 4.3x (Music dataset, budget = 1000) over Almser alone. TransER’s runtimes range from approximately 7 minutes (WDC-computer ) to 2.5 hours (Dexter ), considering the complete training dataset and 50% of it. For datasets consisting of many feature vectors, the clustering approach in MoRER reduces the search space for selecting training data. In contrast to MoRER, TransER compares each unsolved feature vector with all feature vectors from the integrated ER tasks. Ditto is slower than MoRER combined with Almser or Bootstrap due to its training times. In contrast, AnyMatch achieves smaller runtimes by using smaller training datasets. However, neural network-based methods, including Ditto and AnyMatch, generally require more training time than non-language model-based approaches such as MoRER, Almser, and TransER. Notably, Sudowoodo exhibits the longest runtimes, reaching up to 16 hours on the Dexter dataset. As expected, MultiEM outperforms the supervised ML methods in terms of runtimes due to the missing training data and model generation step.
Table 4 presents a summary of the speedup factors for MoRER in comparison to other supervised ML methods. The analysis highlights that using Bootstrap as an AL method within MoRER results in significant runtime improvements over Almser, especially when compared to Almser alone, TransER, Ditto, and Sudowoodo. In terms of effectiveness, MoRER combined with the Bootstrap AL method leads to a decrease in quality compared to the Almser combination. Due to the missing training data generation step, the supervised-based MoRER achieves the best speedup compared to the other approaches (not considered in Table 4).
@@ 5.4. Analysis of Distribution Tests
This section compares the various distribution tests for computing the similarity between ER problems. The similarity is essential for grouping tasks and determining which model $M_{C}^{i}$ of a cluster $C^{i}_{P}$ is used for a new ER problem $D_{l}\in\mathcal{N}$. Fig. 7 presents the results of the three distribution tests: Kolmogorov-Smirnov (KS), Wasserstein distance (WD), and Population Stability Index (PSI). We consider the different active learning methods.
@@
The Kolmogorov-Smirnov test, on the other hand, exhibited strong overall performance and proved to be a reliable option across various datasets and AL methods. It slightly outperformed the other tests for the Dexter dataset, demonstrating its effectiveness in handling heterogeneous data.
5.5. Amount of solved Entity Resolution Tasks
Table 5. Comparison of different ratios of initial ER problems $\mathcal{P_{I}}$ considering the Dexter dataset.
Almser Bootstrap
Budget $ratio_{init}$ F1 $std_{F1}$ F1 $std_{F1}$
1000 30% 0.83 0.067 0.9 0.029
50% 0.934 0.001 0.89 0.012
1500 30% 0.939 0.003 0.79 0.015
50% 0.94 0.001 0.89 0.024
2000 30% 0.84 0.029 0.895 0.017
50% 0.93 0.001 0.9 0.017
We conclude by evaluating the effectiveness of our method based on the proportion of solved ER tasks, reducing the ratio from 50% to 30%. Due to the limited number of ER problems in the WDC-computer and Music datasets, we focus solely on the Dexter dataset.
The proportion of solved ER tasks directly impacts both the number of clusters and the number of tasks within each cluster. As a result, the candidate space for selecting informative similarity feature vectors is also reduced. Additionally, increasing unsolved tasks may cause a growing problem shift, resulting in less effective cluster models for new ER tasks. Table 5 presents the outcomes of reducing the number of solved ER tasks. The results demonstrate that, despite the reduction, the Bootstrap AL method achieves comparable linkage quality with budgets of 1,000 and 2,000. However, when combined with Almser, similar performance is only achieved with a budget of 1,500 compared to a 50% ER problem ratio. This suggests that selecting informative links from a smaller candidate set due to smaller clusters is more challenging, as indicated by the higher standard deviation for the $ratio_{init}=30\%$.
@@ 5.6. Analysis of selection methods
The reclustering selection strategy leads to improved linkage quality compared to the base selection strategy. The $sel_{\text{cov}}$ achieves the best results for the WDC-computer and Music datasets, utilizing a threshold $t_{cov}$ of 0.1. However, these strategies also require the most additional labeling effort compared to the other selection methods. In the Dexter dataset, we observe a negative impact from a small covered ratio, such as 0.1. Due to the wide range of sizes of ER problems in the Dexter dataset, a single ER problem can trigger unnecessary retraining, as indicated by higher F1 scores when using a threshold $t_{cov}$ greater than 0.1.
@@ 6. Discussion
RQ2: Which active learning approach should be used? As the results indicate, our modified Bootstrapping outperforms the baseline methods and the combination with Almser regarding efficiency and achieves comparable results for clean ER problems. However, for dirty ER problems where data sources with duplicates are involved and a high number of problems, the combination with Almser should be used.
In summary, Almser benefits from the reduced search space of candidate record pairs, allowing Almser to focus on the most informative instances for labeling. The combination can be applied to heterogeneous and noisy datasets. If runtime efficiency is a critical requirement and the datasets are cleaner (e.g., Music dataset), the Bootstrap AL method can be considered.
RQ3: Which distribution test for computing the similarity between ER problems is the most effective? The Wasserstein distance results in the lowest F1 score for the bootstrapping-based AL approach on the Dexter and Music datasets. We hypothesize that the Wasserstein distance generates overly similar results for these datasets, which negatively affects performance, particularly in datasets with many ER tasks (Dexter) or a large number of feature vectors (Music). However, the Wasserstein distance outperforms the Kolmogorov-Smirnov test for the impure and smaller WDC-computer dataset, utilizing a certain budget. Meanwhile, the Population Stability Index achieves moderate linkage quality across the datasets for the Bootstrap AL method.
@@
RQ4: What are the limitations of our method? Our method assumes the availability of a sufficient number of solved ER problems. However, when the number of solved problems is limited, the diversity of feature vectors becomes constrained. This limitation can result in models that fail to generalize effectively across the broader spectrum of ER problems, leading to a significant domain shift between solved and unsolved problems. To address this issue, we proposed selection strategies integrating the new ER problems in combination with an updating mechanism of the existing models. This process ensures that the models remain adaptable to evolving data distributions. In this work, we only consider the heterogeneity regarding the feature space. Therefore, we can only apply our method to ER problems with common features. A trivial strategy to handle data sources with different attributes is to apply a pre-trained language model to determine a comparable record representation.
@@ 7. Conclusion
The evaluation demonstrates comparable quality to Almser and superior results to TransER, with significant runtime improvements of up to 40 times faster than Almser and 102 times faster than TransER, highlighting the method’s scalability and practical value. Compared to the language model-based approaches, MoRER achieves comparable performance or outperforms Ditto using 50% of the training data, ranging from roughly 23K to 115K labeled record pairs. Comparing MoRER with the small language model-based approach AnyMatch, Sudowoodo, and MultiEM, our method significantly outperforms them.
In future work, we aim to integrate our method into a comprehensive system for storing and querying solved ER problems, clusters, and classification models, enabling solution reuse for new tasks. While our approach assumes a shared feature space, we suggested the combination of attributes into a single embedding using pre-trained language models. We also plan to explore incrementally expanding the feature space for specific ER tasks.
@@
The reference code is available from our repository at https://github.com/vicolinho/Model_Reuse_ER. The used entity resolution tasks and datasets are available at https://cloud.scadsai.uni-leipzig.de/index.php/s/G64bf9YEy8xYYbK.

### v3

@@ D B MoRER+Almser MoRER+BS Almser Sudowoodo AnyMatch B MoRER Ditto Unicorn TransER
Ditto (Li et al., 2020) and Unicorn (Fan et al., 2024). For faster training, we utilize DistilBERT (Sanh et al., 2019) as a pre-trained model. Moreover, for Ditto, we use a maximum sequence length of 256, also fixed in the original work (Li et al., 2020), and train the model for 10 epochs. We report the results that lead to the smallest training loss over the 10 epochs. For Unicorn, we use the default parameter values of the implementation on GitHub [^6] https://github.com/ruc-datalab/Unicorn with six experts. We utilize the combined loss function covering the diversity and a balanced importance of experts.
We evaluate AnyMatch (Zhang et al., 2025) using its official GitHub implementation [^7] https://github.com/Jantory/anymatch. For each dataset, the model is trained on the corresponding training set and evaluated on the test set, with the parameterized sample size $n_{r}$ for comparability. Note, AnyMatch filters relevant record pairs using the full ground truth of the training data. We use a batch size of 32, GPT-2 as the language model, and a learning rate of $2^{-5}$.
5.3. Comparison to Baselines
Effectiveness. The results, shown in Table 4, indicate that MoRER performs comparably to Almser under certain budgets and active learning configurations. MoRER, when combined with Almser, outperformed both the Bootstrap method and Almser alone on the Dexter dataset. We hypothesize that Almser benefits from the high number of entity resolution tasks, resulting in rich similarity graphs, even when these graphs are divided due to the separation of ER tasks. Moreover, the partitioning of ER tasks reduces the search space for identifying informative links for similar classes of entity resolution tasks, thereby minimizing the likelihood of selecting uninformative feature vectors.
In contrast, TransER consistently underperformed relative to the combined performance of Almser and MoRER across all tasks, regardless of whether the full training dataset or the 50% version was utilized. This consistent underperformance shows its limitations in handling large volumes of heterogeneous training data, which is particularly evident in the Dexter and WDC-computer datasets. Considering the language model-based methods without using labeled training data, Sudowoodo and MultiEM achieve comparable results compared to the aforementioned methods according to the Music dataset. However, the approaches result in a decreased F1 score of up to 38% (Dexter) when comparing Sudowoodo with MoRER+Almser. Sudowoodo was originally conceptualized for pairwise record linkage. Therefore, we assume that the various and heterogeneous ER problems in Dexter and WDC-computer cause the challenge of determining representations to distinguish matches from non-matches.
Table 5. Summary of speedup factors for MoRER with various AL methods compared to the approaches (Alm)ser, TransER (TER), (Su)dowoodo, (Dit)to, (Uni)corn, and (Any)Match regarding the datasets (Dex)ter, (Mus)ic, and (WDC)-computer.
@@ 5.3. Comparison to Baselines
Dex 27.3 127.3 102.7 46 413.8 410.5 45.1 65 132.2 28.1 96.6 78.0 41.2 314.1 311.6 34.3 49.4 106.1 28.9 72.8 58.7 33.4 236.5 234.6 25.8 37.2 76.8
Mus 75.7 33.1 40.6 80.6 87.4 115.4 49 92.8 74.3 21.9 20.2 24.8 49.9 53.5 70.6 30 56.7 14 63.8 15.3 18.8 40 40.4 53.4 22.7 42.9 11.2
WDC 9.8 2.3 2.4 306 16.0 21.0 6.8 12.8 17.3 13.2 1.8 1.9 240.5 12.5 16.4 5.3 10 15.4 15.0 1.3 1.4 176.7 8.9 11.8 3.8 7.2 11.2
AnyMatch and Unicorn achieve comparable performance only on the Music dataset. For the WDC-computer dataset, the results of AnyMatch decrease by up to 22% compared to MoRER+Almser and by 17.4% when using MoRER with Bootstrap. We assume this decrease is caused by the limitations of the AutoML selection, which appears insufficient for building effective models when classifying large numbers of record pairs. For Dexter and WDC-computer, Unicorn generates results that decreased by up to 25% in F1 score compared to MoRER+supervised.
@@
Efficiency. We consider the total runtime that encompasses training data generation, model training, and classification. Notably, for AL methods, the time required to select training data significantly impacts the overall labeling process, since the selection of informative samples often is complex and computationally intensive. MoRER combined with the Bootstrap AL method demonstrates significant advantages, outperforming both Almser and TransER across all datasets and labeling budgets, as shown in  Fig. 5. The additional effort required for statistical analysis, ER clustering and model selection is minimal, ranging from 6 seconds (Music) to 121 seconds (WDC-computer). Compared to selecting training data using an AL method, the proportion ranges from 0.3% (Music with Almser ) to 67% (Dexter with Bootstrap).
For example, on the Music dataset, MoRER+Bootstrap runtimes range from 121 to 2621 seconds, whereas Almser can take up to 4.6 hours. Almser’s longer runtimes are due to the increasing size of similarity graphs, which require more computational resources for graph analysis. The efficiency advantage results from the MoRER strategy of clustering similar ER tasks among the integrated data sources to select training data. This clustering approach significantly decreases the search space for potential links, contributing to the observed performance improvements. Reducing the search space also positively influences the combination of MoRER and Almser. The performance boost results in a maximum speedup of 4.3x (Music dataset, budget = 1000) over Almser alone. TransER’s runtimes range from approximately 7 minutes (WDC-computer) to 2.5 hours (Dexter), considering the complete training dataset and 50% of it.
For datasets consisting of many feature vectors, the clustering approach in MoRER reduces the search space for selecting training data. In contrast to MoRER, TransER compares each unsolved feature vector with all feature vectors from the integrated ER tasks. Ditto is slower than MoRER combined with Almser or Bootstrap due to its training times. In contrast, AnyMatch and Unicorn achieve smaller runtimes by using smaller training datasets resp. less complex models. However, neural network-based methods, including Ditto, Unicorn, and AnyMatch, generally require more training time than non-language model-based approaches such as MoRER, Almser, and TransER. Notably, Sudowoodo exhibits the longest runtimes, reaching up to 16 hours on the Dexter dataset.
Table 5 presents a summary of the speedup factors for MoRER in comparison to other supervised ML methods. The analysis highlights that using Bootstrap as an AL method within MoRER results in significant runtime improvements over Almser, especially when compared to Almser alone, TransER, Ditto, and Sudowoodo. In terms of effectiveness, MoRER combined with the Bootstrap AL method leads to a decrease in quality compared to the Almser combination. Due to the missing training data generation step, the supervised-based MoRER achieves the best speedup compared to the other approaches.
@@ 5.4. Analysis of Distribution Tests
This section compares the various distribution tests for computing the similarity between ER problems. The similarity is essential for grouping tasks and determining which model $M_{C}^{i}$ of a cluster $C^{i}_{P}$ is used for a new ER problem $D_{l}\in\mathcal{N}$. Fig. 6 presents the results of the three distribution tests: Kolmogorov-Smirnov (KS), Wasserstein distance (WD), Population Stability Index (PSI), and the classifier two-sample test (C2ST).
@@
The Kolmogorov-Smirnov test and the multivariate analysis classifier two-sample test, on the other hand, exhibited strong overall performance and proved to be reliable options across various datasets and AL methods. It slightly outperformed the other tests for the Dexter dataset, demonstrating its effectiveness in handling heterogeneous data.
@@ 5.5. Analysis of Selection Methods
The reclustering selection strategy leads to improved ER quality compared to the base selection strategy. The $sel_{\text{cov}}$ achieves the best results for the WDC-computer and Music datasets, utilizing a threshold $t_{cov}$ of 0.1. However, these strategies also require the most additional labeling effort compared to the other selection methods. In the Dexter dataset, we observe a negative impact from a small covered ratio, such as 0.1. Due to the wide range of sizes of ER problems in the Dexter dataset, a single ER problem can trigger unnecessary retraining, as indicated by higher F1 scores when using a threshold $t_{cov}$ greater than 0.1.
@@ 6. Discussion
RQ2: What are the additional costs of our method? To reuse models for new ER tasks, we initially compute pairwise the similarities between the available ER tasks and cluster the resulting ER problem similarity graph. The similarity computation and the clustering scale with the number of initial ER tasks. However, the proportion of the total runtime is below 1% with Almser and 70% using Bootstrap. Moreover, due to the reduced sets of record pairs for each cluster, the selection of informative record pairs using an AL method is more efficient than the application on the union of all ER tasks. In a real-world application, we would apply the retraining of clusters described in Sect. 4.5 due to potential domain shifts. Consequently, we would run the clustering step for each new ER tasks, which is a moderate effort due to the efficiency of the Leiden algorithm (Traag et al., 2019). Moreover, we would retrain the underlying models with new selected training data from the upcoming ER tasks.
RQ3: Which active learning approach should be used? As the results indicate, our modified Bootstrapping outperforms the baseline methods and the combination with Almser regarding efficiency and achieves comparable results for clean ER problems. However, for dirty ER problems where data sources with duplicates are involved and a high number of problems, the combination with Almser should be used. Almser benefits from the reduced search space of candidate record pairs, allowing Almser to focus on the most informative instances for labeling. The combination can be applied to heterogeneous and noisy datasets. If runtime efficiency is a critical requirement and the datasets are cleaner (e.g., Music dataset), the Bootstrap AL method can be considered.
RQ4: Which distribution test for computing the similarity between ER problems is the most effective? The Wasserstein distance results in the lowest F1 score for the bootstrapping-based AL approach on the Dexter and Music datasets. We hypothesize that the Wasserstein distance generates overly similar results for these datasets, which negatively affects performance, particularly in datasets with many ER tasks (Dexter) or a large number of feature vectors (Music). However, the Wasserstein distance outperforms the Kolmogorov-Smirnov test for the impure and smaller WDC-computer dataset, utilizing a certain budget. Meanwhile, the Population Stability Index achieves moderate linkage quality across the datasets for the Bootstrap AL method.
@@
RQ5: What are the limitations of our method? Our method assumes the availability of a sufficient number of solved ER problems. However, when the number of solved problems is limited, the diversity of feature vectors becomes constrained. This limitation can result in models that fail to generalize effectively across the broader spectrum of ER problems, leading to a significant domain shift between solved and unsolved problems.
To address this issue, we proposed selection strategies integrating the new ER problems in combination with an updating mechanism of the existing models. This process ensures that the models remain adaptable to evolving data distributions. In this work, we only consider the heterogeneity regarding the feature space. Therefore, we can only apply our method to ER problems with common features. A trivial strategy to handle data sources with different attributes is to apply a pre-trained language model to determine a comparable representation.
@@ 7. Conclusion
The evaluation demonstrates comparable quality to Almser and superior results to TransER, with significant runtime improvements, highlighting the method’s scalability and practical value. Compared to the language model-based approaches, MoRER achieves comparable performance or outperforms Ditto using 50% of the training data, ranging from roughly 23K to 115K labeled record pairs. Comparing MoRER with the language model-based approaches AnyMatch, Unicorn, and Sudowoodo, our method significantly outperforms them.
In future work, we aim to integrate our method into a comprehensive system for storing and querying solved ER problems, clusters, and classification models, enabling solution reuse for new tasks. While our approach assumes a shared feature space, we suggested the combination of attributes into a single embedding using pre-trained language models. We also plan to explore incrementally expanding the feature space for specific ER tasks. The effectiveness of the reuse depends on the clusters based on the ER problem graph and the similarities. As an additional criterion for the model performance of a cluster, we will investigate the relationship between model performance and cluster stability measures.
@@
The reference code is available from our repository at https://github.com/vicolinho/Model_Reuse_ER.
@@ Acknowledgment
This work was partially funded by Universities Australia and the German Academic Exchange Service (DAAD) grant 57701258.